3) An aluminum cup contains 225 g of water and a 40-g copper stirrer, all at 27°C. A 487-g sample of silver at an initial temperature of 86°C is placed in the water. The stirrer is used to stir the mixture gently until it reaches its final equilibrium temperature of 32°C. Calculate the mass of the aluminum cup. ______g
Answer #1

formula heat Q=mc(T2-T1)

The heat given off by the silver is 0.234*487*54 = 6153.732 J

where temp, 54C.= 86-32°
0.234 J/g*K = specific heat capacity c of Ag

the heat taken up by Al, Cu, H20 is
heat = 0.896*x*5 + 0.381*40*5 + 4.18*225*5

where 0.896, 0.381, and 4.18 = sp. heat cap. of Al, Cu, and H2O
and T= 5° = 32-27°

set both energies equal:
6153.732 = 0.896*x*5 + 0.381*40*5 + 4.18*225*5 J --> solve for x:
x = 306.92g = mass of aluminium cup